# CASS: Nvidia to AMD Transpilation with Data, Models, and Benchmark

### v2

@@ 4.1 Dataset Analysis
Opcode Diversity. We noticed that tensor operations dominate both CUDA and HIP assembly, especially in device code, with memory-related instructions such as mov and call appearing most frequently (refer to appendix A.4). Additionally, HIP opcodes like s_mov_b32 and v_add_co_u32 are used extensively reflecting low-level vector and memory operations unique to AMD’s ISA, while Nvidia is dominated by its own variant of common instructions such as movq, call, and jmp, with greater host-side integration (refer to appendix A.4). Both stacks share common control and memory ops (e.g., mov, test), but HIP provides finer-grained access to GPU internals, revealing deeper visibility into parallelism. The synthetic subset emphasizes memory-oriented instructions, aligning with LLM-driven template optimizations. Figure 6 further shows t-SNE clusters of opcode embeddings (via BERTCoder), suggesting that despite backend differences, Nvidia and AMD share semantically aligned opcode distributions across device and host levels.
@@ A.1 Evaluation on ZLUDA
To assess ZLUDA’s ability to execute CUDA code on AMD GPUs, we designed a two-track evaluation strategy targeting both source-level and binary-level workflows (the latter being akin to assembly-level translation). In the source-to-source setting, we leveraged access to the original CUDA source files to manually compile them into PTX using nvcc. These PTX files were then ingested by ZLUDA, which translated them into AMD-compatible LLVM IR before lowering them into native executables targeting RDNA3 hardware. In the assembly-to-assembly setting, we instead compiled the CUDA source into a complete executable and invoked it directly. ZLUDA intercepted the CUDA runtime calls, dynamically translated the embedded PTX or SASS, and executed the resulting code on the AMD backend. This dual strategy allowed us to assess both ZLUDA’s static translation capabilities and its runtime interoperability under realistic execution conditions.
@@ A.3 CASS Domain Coverage
To obtain the domain-level breakdown shown in Figure 3, we developed a static analysis pipeline that categorizes each source file based on its content. The classification is performed by matching the file’s text against curated sets of domain-specific keywords corresponding to seven high-level categories: general compute, simulation, data structure, machine learning, graphics, cryptography, and scientific computing. Each keyword set includes terms commonly associated with the respective domain; for example, the machine learning category includes terms such as neural, gradient, and activation, while cryptography includes hash, encrypt, and signature. For a given file, the domain with the highest keyword match count is assigned. If no keywords are matched, a default label (e.g., general compute) is applied. After all files are processed, their assignments are aggregated to produce the final domain distribution. This process provides a simple yet straightforward and interpretable way of grouping source files by their functional domain.
@@ A.4.3 Opcode Diversity
Taking a deeper dive into the low-level instructions representation shown in figure 9, a few extra insights can be drawn. In the HIP case, many opcodes, such as s_mov_b32, v_add_co_u32, and s_waitcnt, come directly from AMD’s GPU instruction set. These reflect fine-grained control over the hardware, including scalar and vector operations and synchronization. On the other hand, the CUDA assembly is mostly made up of x86-64 instructions like movq, call, jmp, and pushq, which are typically used on the CPU. This suggests that the CUDA output includes more host-side code or that GPU instructions are hidden behind a higher level of abstraction. Still, both stacks share common instructions like mov and test, showing that some basic control and memory operations are similar. In general, HIP provides more visibility into what the GPU is doing, while CUDA hides many of those low-level details behind a more unified host-device model.
@@ A.5.2 Qualitative Comparison with Other LLMs
We highlight several cases where CASS-7B outperforms existing LLMs such as Claude, Qwen-Coder, and GPT-4o in faithfully transpiling CUDA to HIP. For example, in one instance, CASS-7B correctly transpiled the CUDA code while preserving the exact string constants from the original program, including the label CUDA in the output format string. Maintaining these strings is essential for preserving the intended user-facing behavior, particularly in logging or debugging scenarios where clarity and consistency matter. In contrast, Claude, Qwen-Coder, and GPT4o unnecessarily altered the string to say HIP, despite the output still originating from a CUDA kernel. This substitution introduces a semantic error, as the original string refers to CUDA, not HIP, and should remain unchanged.
@@
In another example, CASS-7B retained the classical CUDA-style kernel launch syntax using triple angle brackets (<<<...>>>), while also ensuring that the generated code remained compilable by correctly including the required HIP header <hip/hip_runtime.h>. This demonstrates a high degree of structural fidelity to the source code, which is especially important for developers familiar with standard CUDA conventions. In contrast, other models such as Claude and Qwen-Coder replaced the launch expression with the HIP-specific macro hipLaunchKernelGGL, which, while functionally valid, deviates from the original representation. More critically, they failed to include the necessary HIP header, rendering the output uncompilable. This example highlights how CASS-7B goes beyond syntactic accuracy to produce code that is both faithful to the original structure and immediately usable in a real compilation setting.

### v3

@@ 4.1 Dataset Analysis
Opcode Diversity. We noticed that tensor operations dominate both CUDA and HIP assembly, especially in device code, with memory-related instructions such as mov
and call
appearing most frequently (refer to appendix A.4). Additionally, HIP opcodes like s_mov_b32
and v_add_co_u32
are used extensively reflecting low-level vector and memory operations unique to AMD’s ISA, while Nvidia is dominated by its own variant of common instructions such as movq
, call
, and jmp
, with greater host-side integration (refer to appendix A.4). Both stacks share common control and memory ops (e.g., mov
, test
), but HIP provides finer-grained access to GPU internals, revealing deeper visibility into parallelism. The synthetic subset emphasizes memory-oriented instructions, aligning with LLM-driven template optimizations. Figure 6 further shows t-SNE clusters of opcode embeddings (via BERTCoder), suggesting that despite backend differences, Nvidia and AMD share semantically aligned opcode distributions across device and host levels.
@@ A.1 Evaluation on ZLUDA
To assess ZLUDA’s ability to execute CUDA code on AMD GPUs, we designed a two-track evaluation strategy targeting both source-level and binary-level workflows (the latter being akin to assembly-level translation). In the source-to-source setting, we leveraged access to the original CUDA source files to manually compile them into PTX using nvcc
. These PTX files were then ingested by ZLUDA, which translated them into AMD-compatible LLVM IR before lowering them into native executables targeting RDNA3 hardware. In the assembly-to-assembly setting, we instead compiled the CUDA source into a complete executable and invoked it directly. ZLUDA intercepted the CUDA runtime calls, dynamically translated the embedded PTX or SASS, and executed the resulting code on the AMD backend. This dual strategy allowed us to assess both ZLUDA’s static translation capabilities and its runtime interoperability under realistic execution conditions.
@@ A.3 CASS Domain Coverage
To obtain the domain-level breakdown shown in Figure 3, we developed a static analysis pipeline that categorizes each source file based on its content. The classification is performed by matching the file’s text against curated sets of domain-specific keywords corresponding to seven high-level categories: general compute, simulation, data structure, machine learning, graphics, cryptography, and scientific computing. Each keyword set includes terms commonly associated with the respective domain; for example, the machine learning category includes terms such as neural
, gradient
, and activation
, while cryptography includes hash
, encrypt
, and signature
. For a given file, the domain with the highest keyword match count is assigned. If no keywords are matched, a default label (e.g., general compute) is applied. After all files are processed, their assignments are aggregated to produce the final domain distribution. This process provides a simple yet straightforward and interpretable way of grouping source files by their functional domain.
@@ A.4.3 Opcode Diversity
Taking a deeper dive into the low-level instructions representation shown in figure 9, a few extra insights can be drawn. In the HIP case, many opcodes, such as s_mov_b32
, v_add_co_u32
, and s_waitcnt
, come directly from AMD’s GPU instruction set. These reflect fine-grained control over the hardware, including scalar and vector operations and synchronization. On the other hand, the CUDA assembly is mostly made up of x86-64 instructions like movq
, call
, jmp
, and pushq
, which are typically used on the CPU. This suggests that the CUDA output includes more host-side code or that GPU instructions are hidden behind a higher level of abstraction. Still, both stacks share common instructions like mov
and test
, showing that some basic control and memory operations are similar. In general, HIP provides more visibility into what the GPU is doing, while CUDA hides many of those low-level details behind a more unified host-device model.
@@ A.5.2 Qualitative Comparison with Other LLMs
We highlight several cases where CASS-7B outperforms existing LLMs such as Claude, Qwen-Coder, and GPT-4o in faithfully transpiling CUDA to HIP. For example, in one instance, CASS-7B correctly transpiled the CUDA code while preserving the exact string constants from the original program, including the label CUDA
in the output format string. Maintaining these strings is essential for preserving the intended user-facing behavior, particularly in logging or debugging scenarios where clarity and consistency matter. In contrast, Claude, Qwen-Coder, and GPT4o unnecessarily altered the string to say HIP
, despite the output still originating from a CUDA kernel. This substitution introduces a semantic error, as the original string refers to CUDA, not HIP, and should remain unchanged.
@@
In another example, CASS-7B retained the classical CUDA-style kernel launch syntax using triple angle brackets (<<<...>>>
), while also ensuring that the generated code remained compilable by correctly including the required HIP header <hip/hip_runtime.h>. This demonstrates a high degree of structural fidelity to the source code, which is especially important for developers familiar with standard CUDA conventions. In contrast, other models such as Claude and Qwen-Coder replaced the launch expression with the HIP-specific macro hipLaunchKernelGGL, which, while functionally valid, deviates from the original representation. More critically, they failed to include the necessary HIP header, rendering the output uncompilable. This example highlights how CASS-7B goes beyond syntactic accuracy to produce code that is both faithful to the original structure and immediately usable in a real compilation setting.